# The Compton-Getting origin of the large-scale anisotropy of Galactic cosmic rays

### v2

@@ 2.4 Compton-Getting effect
Furthermore, the motion of the Solar System through the Galaxy introduces an energy-independent dipole anisotropy due to the Compton-Getting effect (Compton & Getting, 1935):
$\boldsymbol{\delta}=(2+\gamma)\dfrac{\boldsymbol{v}_{\odot}}{c}~{},$
where $v_{\odot}$ is the velocity of the Sun and $\gamma$ is the CR spectral index (Forman, 1970). The velocity is initially thought to be the circular speed of the Sun around the Galactic center, which is approximately $220$ km/s. The dipole anisotropy is expected to be $2.3\times 10^{-3}$ with its peak at $\alpha\simeq 315^{\circ},\delta\simeq 0^{\circ}$. However, this expectation is inconsistent with YBJ experiment results at $\sim 300$ TeV, suggesting that CRs co-rotate with the average stellar motion.
However, the exact rest frame of the cosmic rays is ambiguous. Moreover, many observations have shown that the Sun is not stationary relative to the average motion of the local stars and the local interstellar medium, and its relative speed is smaller than $220$ km/s. If the overall cosmic rays remain at rest with the average motion of the local stars or the local interstellar medium, the relative motion of the Sun could induce a dipole anisotropy. Two kinds of candidates can be considered for the local rest reference frame. One is the so-called local standard of rest (LSR). It is one situation in space with a velocity equal to the average velocity of stars in the solar neighborhood, including the Sun. By the local standard of rest, the space velocities of stars average out to zero. The LSR is moving around the center of the Galaxy at a speed of approximately $200\sim 240$ km/s. The motion of the Sun relative to the LSR is found to be $v_{\rm LSR}\approx 18.0\pm 0.9~{}\rm km/s$, which points to $l\approx 47.9^{\circ}\pm 2.9^{\circ}$ and $b\approx 23.8^{\circ}\pm 2.0^{\circ}$ (Schönrich et al., 2010). Another possibility would be a relative velocity through the local interstellar medium (ISM). The velocity vector measured by the IBEX Experiment is $v_{\rm ISM}\approx 23.2\pm 0.3~{}\rm km/s$ with orientation $l\approx 5.25^{\circ}\pm 0.24^{\circ}$ and $b\approx 12.0^{\circ}\pm 0.5^{\circ}$ with (McComas et al., 2012).
@@ 3 Results and Discussion
First, the propagation parameters in the SDP model are determined by fitting the boron-to-carbon (B/C) ratio. The unknown propagation parameters include $D_{0}$, $\delta_{0}$, $N_{m}$, $\xi$, $n$, $v_{A}$, and $z_{h}$. The fitted B/C ratio is shown in Fig. 1, and the corresponding propagation parameter values are listed in Tab. 1. Additionally, the injection spectra of the background nuclei species $i$ are characterized by the normalization $A^{\rm i}$, power-law index $\nu^{\rm i}$, and cut-off rigidity ${\cal R}c$. For local nuclei, the corresponding parameters are $q^{\rm i}_{0}$, $\alpha^{\rm i}$, and ${\cal R}^{\prime}_{c}$. Fig. 2 presents the fitted energy spectra of protons (left) and helium nuclei (right), where the blue, red, and black lines represent contributions from background sources, the Geminga SNR, and their sum, respectively. The injection parameters are summarized in Tab. 2. In the SDP model, high-energy CRs propagate to Earth through the inner halo, while low-energy CRs primarily diffuse in the outer halo. As a result, high-energy CRs arriving at Earth exhibit a harder spectrum, leading to the concave shape observed in the background proton and helium spectra. To account for the observed excesses of both proton and helium fluxes, the injection spectra from the local source must be harder than the averaged background. The cutoff rigidities of different compositions are regarded as the limits of acceleration in the sources and are assumed to be $Z$-dependent. To account for the softening at tens of TeV in proton and helium spectra, the cut-off rigidity of CRs from the Geminga SNR is $30$ TV.
Fig. 1: Comparison between model calculations and observations for B/C ratio. The data points are adopted from AMS-02, CALET and DAMPE measurements (Aguilar et al., 2016; Adriani et al., 2022; Dampe Collaboration, 2022).
@@
      $D_{0}$ [$\mathrm{cm}^{2}\mathrm{s}^{-1}$]       $4.80\times 10^{28}$
@@
      $v_{A}$ [$\mathrm{km}\cdot\mathrm{s}^{-1}$]       6
@@
      $[({\rm m}^{2}\cdot{\rm sr}\cdot{\rm s}\cdot{\rm GeV})^{-1}]$       [PV]       [GeV-1]       [TV]
@@
Fig. 2: Energy spectra of protons (left) and helium nuclei (right). The data points are taken from DAMPE(An et al., 2019; Alemanno et al., 2021), AMS-02 (Aguilar et al., 2015, 2017), CREAM-III (Yoon et al., 2017) respectively. The blue lines are the background fluxes, and the red lines are the fluxes from a nearby Geminga SNR source respectively. The black lines represent the total fluxes.
@@
In Fig. 3, we compare the energy dependence of dipole anisotropy with and without the Compton-Getting (CG) effect. The left panel shows the dipole amplitude: the blue and red solid lines correspond to the expected anisotropies from background sources and the Geminga SNR, respectively. In the SDP model, diffusion in the inner halo is significantly slower than in conventional models, resulting in a much smaller dipole amplitude from background sources below 100 TeV. However, the growth of anisotropy with energy is consistent with observations above 100 TeV. The anisotropy from the Geminga SNR is energy-independent due to time-dependent propagation effects. When the contributions are combined, a bump-like structure appears between 1 TeV and 100 TeV, caused by partial cancellation of CR streaming from background sources and the nearby Geminga SNR, as illustrated by the black dashed line. At around 100 TeV, a local minimum in the dipole amplitude is reached. Beyond this energy, the contribution from the Geminga SNR rapidly diminishes, and background sources dominate the anisotropy.
@@
When the CG effect is included, the anisotropy is further modified. The green solid line in Fig. 3 represents the anisotropy from the CG effect alone. In the local standard of rest (LSR), the Sun’s velocity is approximately 18 km/s, corresponding to a CG amplitude of about $3\times 10^{-4}$, which is energy-independent in both amplitude and phase. The total anisotropy, shown by the black solid line, includes both CR propagation and the CG effect. Since the Sun’s motion direction in the LSR is closer to the Galactic center, the CG effect suppresses the total dipole amplitude between 1 TeV and 100 TeV. More importantly, below $\sim$200 GeV, where anisotropies from both background and local sources diminish to $\sim 10^{-4}$, the CG effect becomes more significant, causing the total anisotropy to increase again at lower energies. This effect can be tested in future experiments.
The CG effect also influences the dipole phase. Because the Sun’s motion in the LSR is oriented toward the Galactic center, the overall dipole phase is shifted down accordingly, as shown by the black solid line in the right panel of Fig. 3. Notably, as energy decreases from several TeV, the dipole phase gradually shifts toward the direction of CG effect. This transition can be tested by current observatories, such as LHAASO and HAWC.
Fig. 3: The model calculations of the amplitude (left) and phase (right) of the dipole anisotropies for proton plus Helium. In the left panel, the blue solid line is the contribution from background sources, the red one is for Geminga SNR, and the cyan one is the result of Compton-Getting effect (CG effect). In the right panel, the cyan and magenta lines represent the directions from CG effect and the Galactic center (GC) respectively. The black solid and dashed lines in both panels are the total anisotropies with and without CG effect. The data points are taken from Milagro (2009; Abdo et al. (2009)), K-Grande (2019; Apel et al. (2019)), IceCube (2010, 2012; Abbasi et al. (2010, 2012)), IceTop (2013; Aartsen et al. (2013)), ARGO-YBJ (2015; Bartoli et al. (2015)), Tibet (2005, 2017; Amenomori et al. (2005, 2017) respectively.
However, current measurements of the Sun’s local velocity still have uncertainties. Fig. 4 compares the CG effects in which the solar motion is relative to the LSR and the local ISM respectively. Their impact on anisotropy is similar despite that they are different measurements, except that the expected total anisotropy below $100$ GeV is larger for the case when solar motion is relative to the local ISM due to the large velocity. Meanwhile, the flip of the dipole phase at lower energy in this case is faster. Fig. 5 illustrates the influence of different measurements of the Sun’s velocity on the dipole anisotropy. As the velocity increases, the total amplitude reduces at TeV energies. However, below TeV, where the contribution from the density gradient of CRs is minimal, the CG effect can introduce substantial changes, when the velocity is large. Additionally, a higher velocity results in a more abrupt and earlier phase transition.
Fig. 4: The model calculations of the amplitude (left) and phase (right) of the dipole anisotropies for the solar motion relative to the LSR (black) and the local ISM (blue) respectively. The solid lines represent the total results and the dashed lines are the contributions from CG effect.
@@
Fig. 5: The amplitude (left) and phase (right) of the dipole anisotropies for different solar velocities, as shown by red, black and blue colors, relative to the LSR frame. Here, the solid lines of different colors are the total anisotropy amplitude and phase, while the dashed lines with the same colors indicate the model calculations of corresponding CG effect.
@@ 4 Summary
In this study, we further investigate the impact of the Sun’s relative motion on the anisotropy, i.e. so-called Compton-Getting effect. The expected amplitude and phase from the CG effect is energy independent. Previously, the CG effect is thought to originate from the Solar circular motion around the Galactic center with an average circular speed $v\simeq 220$ km/s. However, the experiments did not measure such large effect on the anisotropy. Thereupon, it is inferred that the Galactic CRs mostly co-rotate with the Galaxy.
Here, we consider the Sun’s motion relative to the local rest of the standard or local plasma frame. We found that when considering this effect, the overall dipole amplitude decreases and the phase slightly deviates from the direction of the local regular magnetic field at tens of TeV energies. Below hundreds of GeV, when the anisotropy from the density gradient of CRs is small, the influence of the relative motion of the Sun becomes more prominent. Below $\sim 200$ GeV, the dipole amplitude even rises again and approaches the value of the Compton-Getting effect. Meanwhile, at several hundreds of GeV, there is another flip in the dipole phase, which points to the direction of the Compton-Getting effect. These features could be confirmed by future anisotropy measurements in the 100 GeV to TeV energy range.

### v3

@@ 2.4 Compton-Getting effect
Furthermore, the motion of the Solar System through the Galaxy introduces an energy-independent dipole anisotropy due to the CG effect (1935PhRv...47..817C):
$\boldsymbol{\Delta}_{\rm CG}=(2+\gamma)\,\dfrac{\boldsymbol{v}_{\odot}}{c}~,$ (9)
where $v_{\odot}$ is the velocity of the Sun with respect to the LSR or ISM frame and correspondingly $\Delta_{\rm CG}$ stands for either $\rm CG_{LSR}$ or $\rm CG_{ISM}$. $\gamma$ is the CR spectral index (1970P&SS...18...25F). In practice, the total dipole anisotropy is obtained as the vector sum of the gradient-driven contribution [Eq. (8)] and the CG term [Eq. (9)]:
$\boldsymbol{\Delta}_{\rm tot}=\boldsymbol{\Delta}_{\rm grad}+\boldsymbol{\Delta}_{\rm CG}.$ (10)
The velocity is initially thought to be the circular speed of the Sun around the Galactic center, which is approximately $220$ km/s. The dipole anisotropy is expected to be $2.3\times 10^{-3}$ with its peak at $\alpha\simeq 315^{\circ},\delta\simeq 0^{\circ}$. However, this expectation is inconsistent with YBJ experiment results at $\sim 300$ TeV, suggesting that CRs co-rotate with the average stellar motion.
Moreover, the exact rest frame of the cosmic rays is ambiguous. Many observations have shown that the Sun is not stationary relative to the average motion of the local stars and the local interstellar medium, and its relative speed is smaller than $220$ km/s. If the overall cosmic rays remain at rest with the average motion of the local stars or the local interstellar medium, the relative motion of the Sun could induce a dipole anisotropy. Two kinds of candidates can be considered for the local rest reference frame. One is the so-called local standard of rest (LSR). It is defined as the reference frame moving with the average velocity of stars in the solar neighborhood, including the Sun, and has no intrinsic spatial extent (2010MNRAS.403.1829S). In practice, observational determinations of the LSR are based on stellar samples in the solar neighbourhood, typically spanning distances from roughly 100 pc to several hundred parsecs from the Sun. (2021A&A...649A...6G; 2024A&A...687A.286Z). This empirical scale thus remains applicable to our analysis. By the local standard of rest, the space velocities of stars average out to zero. The LSR is moving around the center of the Galaxy at a speed of approximately $200\sim 240$ km/s. The motion of the Sun relative to the LSR is found to be $v_{\rm LSR}\approx 18.0\pm 0.9~\rm km/s$, which points to $l\approx 47.9^{\circ}\pm 2.9^{\circ}$ and $b\approx 23.8^{\circ}\pm 2.0^{\circ}$ (2010MNRAS.403.1829S). Another possibility would be a relative velocity through the local interstellar medium (ISM). The velocity vector measured by the IBEX Experiment is $v_{\rm ISM}\approx 23.2\pm 0.3~\rm km/s$ with orientation $l\approx 5.25^{\circ}\pm 0.24^{\circ}$ and $b\approx 12.0^{\circ}\pm 0.5^{\circ}$ (2012Sci...336.1291M). It’s worth noting that the IBEX-inferred flow direction here is obtained indirectly from charge-exchange measurements of interstellar neutral atoms. Therefore, by simultaneously considering the LSR and the IBEX-based LISM reference frames, we aim to make the evaluation of the CG effect as comprehensive and reliable as possible within the current uncertainties. Furthermore, current observations show no compelling evidence for a substantial bulk motion of the Geminga SNR. Therefore, its motion relative to the local stellar population or the surrounding ISM is neglected in our calculations.
@@ 3 Results and Discussion
Fig. 1: Comparison between model calculations and observations for B/C ratio. The data points are adopted from AMS-02, CALET and DAMPE measurements (2016PhRvL.117w1102A; 2022PhRvL.129y1103A; 2022SciBu..67.2162D).
@@
  $D_{0}$ $\mathrm{cm}^{2}\mathrm{s}^{-1}$]   $4.80\times 10^{28}$
@@
  $v_{A}$ $\mathrm{km}\cdot\mathrm{s}^{-1}$]   6
@@
  $({\rm m}^{2}\cdot{\rm sr}\cdot{\rm s}\cdot{\rm GeV})^{-1}]   [$PV]   [GeV-1]   [TV]
@@
We begin by calibrating the transport parameters of the SDP framework using the boron-to-carbon (B/C) ratio. The key quantities constrained in the fitting process include the diffusion normalization and slope $(D_{0},\delta_{0})$, the modulation factors $(N_{m},\xi,n)$, the Alfvén velocity $v_{A}$, and the halo scale height $z_{h}$. The fitted B/C ratio is shown in Fig. 1, and the corresponding propagation parameter values are listed in Tab. 2. The injection spectra of background nuclei of species $i$ are characterized by the normalization $A_{i}$, spectral index $\nu_{i}$, and cutoff rigidity $R_{c}$. For the nearby component, the corresponding parameters are $(q_{0,i},\alpha_{i},R_{c}^{\prime})$. The injection parameters are summarized in Tab. 3. In the SDP model, high-energy CRs propagate to Earth through the inner halo, while low-energy CRs primarily diffuse in the outer halo. As a result, high-energy CRs arriving at Earth exhibit a harder spectrum, leading to the concave shape observed in the background proton and helium spectra. To account for the observed excesses of both proton and helium fluxes, the injection spectra from the local source must be harder than the averaged background. The cutoff rigidities of different compositions are regarded as the limits of acceleration in the sources and are assumed to be $Z$-dependent. To account for the softening at tens of TeV in proton and helium spectra, the cut-off rigidity of CRs from the Geminga SNR is $30$ TV.
To reproduce the spectral hardening observed around 200 GV, the nearby source is assigned a harder injection spectral index than the background component, and such a difference is physically expected. Observationally, the radio synchrotron spectral index $\gamma$ of shell-type SNRs typically falls within the approximate range 0.8-0.8 to 0.2-0.2, according to Green’s catalog (2023ApJS..265...53R). Using the standard relation between the radio index and the electron spectral index $\alpha=-(\gamma-1)/2$, this corresponds to a cosmic-ray electron spectral index $\alpha$ between approximately $1.4$ and $2.6$. Furthermore, by fitting multiwavelength spectra of 35 SNRs with a one-zone model, 2019ApJ...874...50Z found that both the break energy and the low-energy spectral index (typically $\sim 1.3$$2.3$) decrease with increasing SNR age, implying that younger remnants tend to produce softer spectra with higher break energies, whereas older remnants yield harder spectra with lower break energies. The value of $\alpha\sim 2.1$ inferred for Geminga here comfortably lies within this feasible range, while the softer background spectrum likely reflects the averaged contribution of numerous SNRs with diverse intrinsic spectral properties throughout the Galaxy.
Fig. 2: Top panels: modeled energy spectra of cosmic-ray proton (left) and helium nuclei (right). Measurements from DAMPE (2019SciA....5.3793A; 2021PhRvL.126t1102A), AMS-02 (2015PhRvL.114q1103A; 2017PhRvL.119y1101A), and CREAM-III (2017ApJ...839....5Y) are shown for comparison. The blue curves illustrate the contribution from the large-scale Galactic background, while the red curves indicate the additional flux from a local source associated with the Geminga SNR. The black curves display the combined result of the two components. The corresponding reduced chi-squared ($\chi^{2}$/dof) values are also indicated in the panels. Bottom panels: the fitting residuals. The points of different colors represent the residuals with respect to the AMS-02(red), CREAM(green), and DAMPE(blue), respectively.
Fig. 2 shows the calculated proton and helium spectra together with the corresponding fit residuals [^1] Residuals are defined as $(\Phi_{\rm exp}-\Phi_{\rm mod})/\sigma_{\rm exp}$, where $\Phi_{\rm exp}$ is the measured flux, $\Phi_{\rm mod}$ the model prediction, and $\sigma_{\rm exp}$ the experimental uncertainty.. The top panels present the comparison between the model-predicted spectra and the observational data. To evaluate the quality of the fits, we compute the reduced chi-squared values, which are about 10.8 for protons and 1.2 for helium. The model reproduces the helium spectrum well, whereas the agreement for protons is relatively poorer. This discrepancy is largely driven by the deviation between the model prediction and the AMS-02 measurements. To illustrate this, the bottom panels display the fit residuals, which provide a direct visual measure of the agreement between the model and the data. For AMS-02, the residuals of the proton spectrum are significantly larger than those of helium, resulting in a higher reduced chi-squared value for protons. In contrast, the CREAM and DAMPE data extend to higher energies with larger uncertainties, leading to smaller normalized residuals that remain close to zero for both protons and helium. Nevertheless, the residuals of all three experiments fluctuate around zero within $3\sigma$, confirming that the model reproduces the observed spectra without significant systematic bias.
Fig. 3: The model calculations of the amplitude (left) and phase (right) of the dipole anisotropies for the combined proton and helium flux. In the left panel, the blue solid line is the contribution from background sources, the red one is for Geminga SNR, and the cyan one is the result of Compton-Getting effect ($\rm CG_{LSR}$ effect). In the right panel, the cyan and magenta lines represent the directions from $\rm CG_{LSR}$ effect and the Galactic center respectively. The black solid and dashed lines in both panels are the total anisotropies with and without $\rm CG_{LSR}$ effect. Measurements are taken from Milagro (2009; 2009ApJ...698.2121A), K-Grande (2019; 2019ApJ...870...91A), IceCube (2010, 2012; 2010ApJ...718L.194A; 2012ApJ...746...33A), IceTop (2013; 2013ApJ...765...55A), ARGO-YBJ (2015; 2015ApJ...809...90B), Tibet (2005, 2017; 2005ApJ...626L..29A; 2017ApJ...836..153A respectively.
@@
In Fig. 3, we compare the energy dependence of the dipole anisotropy with and without the CG effect, where the CG term corresponds to the $\rm CG_{LSR}$ contribution. The left panel shows the dipole amplitude: the blue and red solid lines correspond to the expected anisotropies from background sources and the Geminga SNR, respectively. In the SDP model, diffusion in the inner halo is significantly slower than in conventional models, resulting in a much smaller dipole amplitude from background sources below 100 TeV. However, the growth of anisotropy with energy is consistent with observations above 100 TeV. The anisotropy from the Geminga SNR is energy-independent due to time-dependent propagation effects. When the contributions are combined, a bump-like structure appears between 1 TeV and 100 TeV, caused by partial cancellation of CR streaming from background sources and the nearby Geminga SNR, as illustrated by the black dashed line. At around 100 TeV, a local minimum in the dipole amplitude is reached. Beyond this energy, the contribution from the Geminga SNR rapidly diminishes, and background sources dominate the anisotropy.
@@
When the CG effect is included, the anisotropy is further modified. The green solid line in Fig. 3 represents the anisotropy from the $\rm CG_{LSR}$ effect alone. In the local standard of rest (LSR), the Sun’s velocity is approximately 18 km/s, corresponding to a $\rm CG_{LSR}$ amplitude of about $3\times 10^{-4}$, which is energy-independent in both amplitude and phase. The total anisotropy, shown by the black solid line, includes both CR propagation and the $\rm CG_{LSR}$ effect. Since the Sun’s motion direction in the LSR is closer to the Galactic center, the $\rm CG_{LSR}$ effect suppresses the total dipole amplitude between 1 TeV and 100 TeV. More importantly, below $\sim$200 GeV, where anisotropies from both background and local sources diminish to $\sim 10^{-4}$, the $\rm CG_{LSR}$ effect becomes more significant, causing the total anisotropy to increase again at lower energies.
The $\rm CG_{LSR}$ effect also influences the dipole phase. Because the Sun’s motion in the LSR is oriented toward the Galactic center, the overall dipole phase is shifted down accordingly, as shown by the black solid line in the right panel of Fig. 3. Notably, as energy decreases from several TeV, the dipole phase gradually shifts toward the direction of $\rm CG_{LSR}$ effect. Quantitatively, the $\rm CG_{LSR}$ contribution modifies the dipole amplitude by approximately $30\%$$50\%$ in the range $10^{2}$$10^{3}\,\mathrm{GeV}$, and introduces a phase shift of a few hours in local sidereal time. Although the resulting change remains smaller than the current dispersion among experimental determinations, the predicted results—with and without the CG contribution—establish clear and testable benchmarks for ongoing and future ground-based observations. Building on recent LHAASO studies (Liu:2025Pr), the experiment is poised to explore cosmic-ray anisotropy down to energies of a few hundred GeV, opening an exciting window into low-energy cosmic-ray behavior. While factors such as angular and energy resolution or solar modulation will need to be carefully accounted for, these challenges also present valuable opportunities to refine detection techniques and modeling frameworks. Successfully observing the predicted features would mark a significant milestone, providing a direct probe of the local cosmic-ray reference frame and deepening our understanding of cosmic-ray propagation in the heliosphere.
Fig. 4: Comparison of predicted dipole anisotropy amplitude (left) and phase (right) under two assumptions for the solar motion: relative to the LSR (black) and relative to the local ISM (blue). Solid curves show the combined anisotropy, while dashed curves illustrate the individual CG contribution in each case.
@@
Fig. 5: Predicted dipole anisotropy amplitude (left) and phase (right) for different assumed solar velocities with respect to the LSR, shown in red, black, and blue. Solid curves represent the total anisotropy including all contributions, while dashed curves display the corresponding $\rm CG_{LSR}$ terms for each velocity.
However, current measurements of the Sun’s local velocity still have uncertainties. Fig. 4 compares the CG effects in which the solar motion is relative to the LSR and the local ISM respectively, as mentioned above $\rm CG_{LSR}$ and $\rm CG_{ISM}$. Their impact on anisotropy is similar despite that they are different measurements, except that the expected total anisotropy below $100$ GeV is larger for the case when solar motion is relative to the local ISM due to the large velocity. Meanwhile, the flip of the dipole phase at lower energy in this case is faster. Fig. 5 illustrates the influence of different measurements of the Sun’s velocity on the dipole anisotropy. As the velocity increases, the total amplitude reduces at TeV energies. However, below TeV, where the contribution from the density gradient of CRs is minimal, the $\rm CG_{LSR}$ effect can introduce substantial changes, when the velocity is large. Additionally, a higher velocity results in a more abrupt and earlier phase transition.
Fig. 6: Predicted amplitude (left) and phase (right) of the dipole anisotropy including the $\mathrm{CG}_{\mathrm{LSR}}$ contribution for two scenarios: $\Delta\delta=0$ (dashed) and $\Delta\delta=0.3$ (solid). The black solid line has the same meaning as in Fig. 3, while the black dashed line represents the result for $\Delta\delta=0$.
To illustrate the influence of $\delta_{\perp}$, defined as $\Delta\delta\equiv\delta_{\perp}-\delta_{\parallel}$, on the dipole anisotropy, we compute the amplitude and phase including the $\mathrm{CG}_{\mathrm{LSR}}$ contribution for two scenarios: $\Delta\delta=0$ and $\Delta\delta=0.3$, as shown in Fig. 6. At energies below $\sim 100$ TeV, the anisotropy is mainly governed by the nearby Geminga source, CRs preferentially propagate along the local regular magnetic field under strongly anisotropic diffusion ($D_{\perp}\ll D_{\parallel}$), so the difference between $\Delta\delta=0$ and 0.3 is minor and both the amplitude and phase remain nearly identical. At higher energies, however, the large-scale Galactic background dominates. For $\Delta\delta=0$, the transport remains highly anisotropic, suppressing the effective flux perpendicular to the magnetic field, which yields a smaller amplitude and a phase offset from the Galactic-center direction. In contrast, when $\Delta\delta=0.3$, the perpendicular diffusion coefficient increases more rapidly with energy, driving the transport toward quasi-isotropy. Consequently, the amplitude rises and the phase gradually shifts toward the Galactic-center direction. This behavior is consistent with the standard anisotropic diffusion tensor picture and with previous studies showing that the energy dependence of $D_{\perp}$ can steer the observed dipole properties at high energies(1999ApJ...520..204G).
@@ 4 Summary
In this study, we further investigate the impact of the Sun’s relative motion on the anisotropy, i.e. so-called Compton-Getting effect. The expected amplitude and phase from the CG effect is energy-independent. Previously, the CG effect is thought to originate from the Solar circular motion around the Galactic center with an average circular speed $v\simeq 220$ km/s. However, the experiments did not measure such large effect on the anisotropy. Thereupon, it is inferred that the Galactic CRs mostly co-rotate with the Galaxy.
Here, we consider the Sun’s motion relative to the local standard rest or local plasma frame. We found that when considering this effect, the overall dipole amplitude decreases and the phase slightly deviates from the direction of the local regular magnetic field at tens of TeV energies. Below hundreds of GeV, when the anisotropy from the density gradient of CRs is small, the influence of the relative motion of the Sun becomes more prominent. Below $\sim 200$ GeV, the dipole amplitude even rises again and approaches the value of the $\rm CG_{LSR}$($\rm CG_{ISM}$) effect. Meanwhile, at several hundreds of GeV, there is another flip in the dipole phase, which points to the direction of the $\rm CG_{LSR}$($\rm CG_{ISM}$) effect. Quantitatively, the $\rm CG_{LSR}$ contribution alters the dipole amplitude by about $30\%$$50\%$ at $10^{2}$$10^{3}\,\mathrm{GeV}$ and induces a phase shift of a few hours in local sidereal time. These features, shown with and without the $\rm CG_{LSR}$($\rm CG_{ISM}$) effect, could be confirmed by future anisotropy measurements in the 100 GeV to TeV energy range.